OpenStudy (anonymous):

I need to know how to solve this please. *The number of calories in a piece of cake is 20 less than 3 times the number of calories of a scoop of ice cream. Together, the cake and ice cream have a total of 460 calories. What is the number of calories for the piece of cake and the scoop of ice cream? A) 340, and 120 calories B) 340, and 220 calories C) 360, and 220 calories D) 360, and 120 calories

OpenStudy (anonymous):

let x be the number of calories in the cake and y the number of calories in the ice cream. you know that x=3y-20 and x+y=460 you express x in one equation with the other one: x=460-y 460 - y= 3y - 20 480=4y y=120 x=460-120=340

OpenStudy (anonymous):

I got it now, thanks a lot